OpenStudy (memyselfandpi):

Pick a two-digit number greater than 25. Rewrite your two-digit number as a difference of two numbers. Show how to use the identity (x − y)^2 = x^2 − 2xy + y^2 to square your number without using a calculator.

OpenStudy (memyselfandpi):

Please help.

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (memyselfandpi):

Thank You :)

OpenStudy (memyselfandpi):

@Mehek14

OpenStudy (michele_laino):

I can choose 30=35-5

OpenStudy (memyselfandpi):

Ok, how can I work this problem out though? That's what I am having trouble with.

OpenStudy (michele_laino):

next, I can write this: \[900 = {30^2} = {\left( {35 - 5} \right)^2} = {35^2} + {5^2} - 2 \times 35 \times 5=...\]

OpenStudy (memyselfandpi):

900 + 25 - 2 x 35 x 5 = 575, if I am supposed to evaluate

OpenStudy (michele_laino):

we have this step: \[\begin{gathered} {30^2} = {\left( {35 - 5} \right)^2} = {35^2} + {5^2} - 2 \times 35 \times 5 = \hfill \\ \hfill \\ = 1225 + 25 - 350 = ...? \hfill \\ \end{gathered} \]

OpenStudy (memyselfandpi):

1225 + 25 -350 = 900.. Which is 30^2

OpenStudy (michele_laino):

that's right!

OpenStudy (memyselfandpi):

Thank you. :)

OpenStudy (michele_laino):

:)